Summary of Science Run: 7-14 May 2003: Sam Ragland dated on 17 May 2002

Hi All,

I observed a few science targets during this run. Weather was good during the first half of this run and later it became cloudy. I lost almost two clear nights due to IOTA control system problems. Listed below is a delay system for the present array configuration:

N35/S15/N10 (A Fixed) configuration:

delay_sys -sname[0] ne25 -station[0][0] 1887.3711 -station[0][1] 1642.3562 -station[0][2] -0.3581 -station[0][3] 2501.9000 -sname[1] se15 -station[1][0] -985.4137 -station[1][1] 1132.4227 -station[1][2] 0.2148 -station[1][3] 1501.1400 -sname[2] ne10 -station[2][0] 754.9485 -station[2][1] 656.9425 -station[2][2] -0.1432 -station[2][3] 1000.7600 -config a -delta[0][0] 0 -delta[0][1] 0 -delta[0][2] 0 -delta[1][0] 0 -delta[1][1] 0 -delta[1][2] 0 -delta[2][0] 0 -delta[2][1] 0 -delta[2][2] 0 -zero[0][0] -165.39062 -zero[1][0] -269.57056 -zero[2][0] -104.34710 -zero[0][1] 0.0 -zero[1][1] 0.0 -zero[2][1] 0.0

The SD offsets for this delay system are [-0.4, -0.7].

A few engineering aspects:

  • Star Acquisition Camera (SA): The performance of the frame grabber and associated OT tools is excellent. I used them during the second half of my run without any difficulty.

  • Side lopes on the Power Spectrum: Occasionally, side lopes are still seen in the power spectrum. In addition, a peak (of low amplitude) is also seen around 200Hz.

  • Control system problems: (1) To start with, the script 'observe_g3.2' was not working. I found that the script was not loading 'source' and 'setenv'. Pete fixed a bug there. We also leant that while 'iota16' and 'VxWorks' are sufficient to run OT, iota17 is also required to run IDL.

  • (2) After recycling power for 'iota16', I had difficulties to startup the control system. I had two problems: (a) wrong clock settings on 'VxWorks' and (b) clients could not memory map. In detail: 'VxWorks' didn't set it's clock at the rebooting time and was reading the factory clock settings ( year 1970!), although 'iota16' had the correct time settings. We expect 'VxWorks' to get time information from 'iota16' in case the network server (which provides more accurate time signal to 'VxWorks') is down. I think, for some reason that did not occur on that night. Finally, it reset to the correct time setting when I left the system idle for about half an hour (hopefully it got time information from the network server). I think, the 'memory mapping problem' is something to do with the loading sequence of various 'VxWorks' processes. We didn't have this problem in the past when correct startup sequence is followed. Pete recommends the following sequence for startup & shutdown.


    1. start up iota16
    2. start up iota17 (if you want to use "observe" program)
    ><=== Under normal circumstances, this is where we start each night.
    3. turn on VME box. this causes VxWorks to boot on all the CPUs and it will try to set the clock using the NTP server.
    4. load the vxworks programs (vxworks-start in iotamc_g3.2 directory)
    5. load the lmtmc program (lmtmc-start in iotamc_g3.2 directory)
    6. load the daq program (daq-start in iotamc_g3.2 directory) Please note that you have a choice here and you COULD actually forget steps 4 and 5. You would do this to test the camera. The consequence is that the IDL program would most likely show an error on its first attempt to map the memory on the VME box. This should only be a temporary problem.
    7. load the IDL "observe" program (observe_g3.2)


    1. quit the IDL program
    2. quit the lmtmc program
    3. quit the daq and vxworks programs (daq-stop vxworks-stop)
    4. shutdown VME cage.

  • Spikes on PICNIC data: In the beginning of my run, I noticed occasional spikes on one night. Possible noise sources are (a) new UPS units (b) Differential power - VxWorks on dirty power while the other IOTA computers on UPS (c) The presence of 'front panel covers' of 'VxWorks', blocking the ventilation (although the room temperature around the CPLD board was sufficient low - 68 deg F). I opened the panel covers, bypassed the new UPS units and the spikes disappeared. I am not sure about the source of those spikes.

  • Shutter problem: As I mentioned in my last run report, occasionally, Fixed beam shutter doesn't close fully and the D2 shutter doesn't open fully.